Journal entry worksheet 2 Record the entry for direct materials costs and variances. Note: Enter debits before credits Transaction General Journal Debit Credit 1
Answer #1

1)

WHERE ;

SP = STANDARD PRICE

AP = ACTUAL PRICE

SQ= STANDARD QUANTITY

AQ= ACTUAL QUANTITY

SH = STANDARD HOURS

AH = ACTUAL HOURS

SR = STANDARD RATE

AR = ACTUAL RATE

Materials Price Variance

$2600

UNFAVOURABLE

Materials Usage Variance

$15422.4

UNFAVOURABLE

direct material price variance = AQ ( SP - AP )

= 5200 sq. ft( $5.1 - $5.6 )

= 5200 sq. ft ( (-)$0.5 )

=(-) $2600 UNFAVOURABLE

direct materials usage variance = SP ( SQ - AQ )

= $5.1 ( 2176sq. ft-5200sq. ft )

= $0.40 ( (-700 pounds ) )

= (-) $15422.4 UNFAVOURABLE

NOTE: STANDARD QUANTITY CALCULATION = 170units * 12.8sq. ft

= 2176sq. ft

total direct materials variance = ( SQ*SP - AQ*AP )

= (2176sq. ft* $5.1 – 5200sq. ft* $5.6)

= ( $11097.6 - $29120 )

= (-) $18022.4 UNFAVOURABLE

a)

DIRECT MATERIAL PRICE VARIANCE JOURNAL

ACCOUNT

DEBIT

CREDIT

RAW MATERIAL INVENTORY

26520

DIRECT MATERIAL PRICE VARIANCE ACCOUNT

2600

ACCOUNTS PAYABLE

29120

TOTAL

29120

29120

b)

Direct materials usage variance journal

Account

Debit

Credit

Work in process inventory

11097.6

Raw materials inventory

26520

Direct materials quantity variance account

15422.4

Total

26520

26520

2)

Labor Rate Variance

$1212.4

UNFAVOURABLE

Labor Efficiency Variance

$910

FAVOURABLE

direct labour rate variance = ( SR - AR ) AH

= ( $13 - $13.70 ) 1732 hrs

= ( (-)$0.7 ) 1732hrs

= (-) $1212.4 UNFAVOURABLE

= direct labour Efficiency variance = ( SH - AH ) SR

= ( 1802hrs - 1732hrs ) $13

= (-) 70hrs * $13

= $910 FAVOURABLE

NOTE: STANDARD HOURS = 10.6 HOURS * 170 UNITS

= 1802 HOURS

total direct labour variance = ( SH*SR - AH*AR )

= ( 18020hrs*$13 - 1732hrs*13.70 )

= ( $23426 - $23728.4 )

= (-) $302.4 UNFAVOURABLE

a)

ACCOUNT DEBIT CREDIT
WORK IN PROCESS INVENTORY 23426
DIRECT LABOUR PRICE VARIANCE ACCOUNT 1212.4
DIRECT LABOUR EFFICIENCY VARIANCE ACCOUNT 910
WAGES PAYABLE 23728.4
TOTAL 24638.4 24638.4